OpenStudy (calculusxy):

If sweatshirts cost $28 and tee shirts cost $15 and William spent $1440 buying a total of 70 sweatshirts and tee shirts for the school store, how many tee shirts did he purchase?

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you have the 2 equations set up?

OpenStudy (calculusxy):

So this is what I have: \[70($28) + 15s = $1440\]

jimthompson5910 (jim_thompson5910):

70 is the number of t-shirts and sweatshirts combined

jimthompson5910 (jim_thompson5910):

x = # of sweatshirts y = # of t shirts x+y = 70

jimthompson5910 (jim_thompson5910):

`sweatshirts cost $28 and tee shirts cost $15 and William spent $1440` 28x + 15y = 1440 is the other equation

OpenStudy (calculusxy):

Oh okay